Project

General

Profile

NDOS Commissioning Meeting 2011/06/20

Attendance: Jon, Enrique, Peter, Kurt, Susan, Jan, Zeb, Athans, Gavin, Rick K, Greg, Mark, Leon,

Agenda:

  1. DCS Status & Plans:
    1. EPICS or ACNET?
    2. Readiness for cooling of some of golden APDs on 1st diblock on Tuesday?
  2. NDOS Status / Status of Detector-Wide APD Inspections (Jon will briefly report based on email from Mat)
  3. DAQ/Electronics Status & Plans
  4. AOB

DCS Status & Plans

  • Athans: meeting last Friday, decided that EPICS solution would be used for this week. Work remains to be done:
    • GUI (Martin and Gennadiy)
    • Low-level issue that Qiming and Ron are working on: fix IOC so that it can set values.
  • We should be ready to begin cooling tomorrow morning.
  • Susan: will these be marked as test runs? Jon: yes

NDOS Status, Hardware Work

  • Mat and Denis have completed the APD inspection on diblock 1. Obviously bad APDs have been removed. They will continue the inspection and hope be done by Wed.
  • APDs removed from diblock 1 will be sent to Jason at CalTech. These will be tested, and if they fail the test, they will be baked.
  • Jon: rough estimate for turn-around? Leon: should be 30-40/day if there aren't major problems

DAQ Software/Electronics Update & Plans

  • Mark: has something changed with timing chain since horn current was changed? Seeing what looks like neutrinos outside of our analysis window. Peter: not to our knowledge.
  • Peter:
    • Leon noticed that we have less headroom in the DCMs these days. Looks like we're hitting ~100% at 12 MB/s, whereas previously we could get ~19 MB/s. So something has changed, possibly the frame size on the buffer nodes, but Ron will need to investigate when he's done working on DCS.
    • Event display and OnMon keep crashing; Peter is investigating. Mark: his student will soon begin working on the ART version of these. Mark offered to look at stack traces.
  • Kurt: possibly will continue working on moving config. to dB
  • Peter: memory leak in DDS. OpenSlice is coming out w/ new version of DDS that the DAQ group will have to deal with.

AOB (All)

  • Mark: NuMI status? Jon: HE running will likely end today, target will be moved tomorrow morning for LE running. There is a leak in what appears to be the outside of the target, so it's not critical, but the NuMI team is proceeding cautiously.